# Drops of volatile binary mixtures on brush-covered substrates

### v1

@@ Abstract
We introduce a mesoscopic hydrodynamic model for drops of binary mixtures of volatile partially wetting liquids on brush-covered substrates. The formulation as gradient dynamics on an underlying free energy functional allows us to systematically account for cross couplings between the six scalar fields needed to describe the two components within the three phases - brush, drop and gas. The energy accounts for concentration- and brush state-dependent capillarity and wettability, miscibility of the two components within drop and brush, and entropy in the gas. Finally, the usage of the model is illustrated by employing numerical time simulations to study the coupled spreading, evaporation, imbibition and partial demixing of a sessile drop on a brush.
@@ 1 Introduction
Overall, the statics and dynamics of (de)wetting of simple nonvolatile liquids on rigid inert solid substrates are rather well understood [1]. However, at present flexible and adaptive substrates, e.g., (hydro)gels and polymer brushes, are of large interest [2, 3, 4, 5, 6]. In general, films and drops of liquids on solid substrates are often modeled employing mesoscopic hydrodynamics in the form of thin-film (or lubrication) models that incorporate capillarity and wettability [7, 8, 9]. Particularly convenient is their formulation in the form of gradient dynamics on underlying energy functionals [10, 11, 12, 5, 6]. Then, consistency between macroscopic and mesoscopic models is ensured by relations between macroscopic interface energies and wetting energies [13, 14, 6].
@@
In the case of polymer brush-covered substrates [15, 16, 17, 18, 19, 20, 21, 22], thin-film models in gradient dynamics formulation [6] incorporate the free energy of the polymer brush, e.g., via an Alexander-de Gennes approach [23, 24] that in the simplest case accounts for mixing energy and entropy of brush polymer and liquid and entropic stretching [25, 26]. As the polymers are grafted to the supporting solid, translational entropy does not contribute. The resulting thin-film models may be employed to investigate the spreading of drops of nonvolatile [26, 6] and volatile [27] liquids on polymer brushes, the stick-slip motion of three-phase contact lines advancing over an initially dry brush [28, 6], as well as sliding drops on such substrates [6]. Note that, although of high interest [25, 29, 19, 22], the dynamic behavior of drops of liquid mixtures and/or their vapors on/above polymer brushes has not yet been modeled by such thin-film models.
@@
However, such mesoscopic hydrodynamics models are already established for binary mixtures of nonvolatile liquids on rigid inert solids [30, 31, 32, 33]. There, a gradient dynamics formulation [31] allows, e.g., to investigate dewetting processes that are triggered by the coupling of fluctuations in film thickness and concentration [32]. Surface activity due to insoluble or soluble surfactants may also be incorporated [34, 14]. Volatile liquids (including vapor diffusion) may be accounted for in thin-film modeling by confining the system into a small gap between two parallel plates as then vapor diffusion is predominantly lateral while concentrations equilibrate fast across the gap [35]. Such a model allows one to describe the full range from mass transfer-limited to diffusion-limited evaporation/condensation. This is applied to droplets spreading on a brush in Ref. [27] where it is shown that even a small lateral vapor leakage may result in the formation of a long-lived stationary halo of macroscopic extension in the brush profile.
Here, we develop a model that allows one to investigate the dynamics of isothermal drops of binary mixtures of volatile partially wetting liquids on a polymer brush within a narrow gap. In particular, we employ the gradient dynamics approach to mesoscopic hydrodynamics. We emphasize that the approach “semi-automatically” guarantees thermodynamic consistency and nevertheless incorporates the relevant dissipation channels.
@@
The model is introduced in section 2 by detailing the six scalar fields representing the amount of the two substances in the three phases brush, drop and gas, and the corresponding dynamic equations. The underlying energy functional and mobility functions are detailed in Appendices A and B. Using the established model, section 3 presents exemplary results obtained by numerical time simulations. Finally, section 3 summarizes our results, points out possible applications of the presented model, and proposes future extensions that may alleviate present limitations.
@@ 2 Mesoscopic hydrodynamic model in gradient dynamics form
We consider an isothermal system as sketched in Fig. 1: A drop of a binary mixture of volatile liquids is located on a swelling polymer brush. Describing the amounts of the two liquid components in the drop and brush by effective heights $\zeta_{\mathrm{d}i}$ and $\zeta_{\mathrm{b}i}$ (with $i=1,2$), respectively, the thickness of the drop is $h_{\text{d}}=\zeta_{\text{d1}}+\zeta_{\text{d2}}$ while the brush thickness is $h_{\text{b}}=\zeta_{\text{b1}}+\zeta_{\text{b2}}+\zeta_{\text{p}}$. Here, $\zeta_{\text{p}}$ is the uniform and constant effective height of polymers, i.e., the dry brush height $\zeta_{\text{p}}=\sigma Nl_{\text{K}}$, with $\sigma$ the relative grafting density, $N$ the number of monomers per polymer chain and $l_{\text{K}}$ the Kuhn length. We consider the system to be positioned inside a narrow gap between two horizontal plates of distance $d$, therefore, the height of the gas phase is $h_{\text{v}}=d-h_{\text{d}}-h_{\text{b}}$.
Thin-film models for liquid films and drops involving mixtures are conveniently written in gradient dynamics form employing (effective) heights [32, 33, 34, 36]. However, when vapor is incorporated, it is preferable to express the amounts of each component in each phase - brush, drop and gas - as particle number densities per substrate area $\psi_{i}(\mathbf{r},t)$, where $\mathbf{r}=(x,y)^{T}$ denotes the substrate coordinates [27, 6]. Here, we use the fields $(\psi_{1},\psi_{2},\psi_{3},\psi_{4},\psi_{5},\psi_{6})=(\rho_{\mathrm{l}1}\zeta_{\text{d1}},\rho_{\mathrm{l}1}\zeta_{\text{b1}},\rho_{\mathrm{v}1}h_{\text{v}},\rho_{\mathrm{l}2}\zeta_{\text{d2}},\rho_{\mathrm{l}2}\zeta_{\text{b2}},\rho_{\mathrm{v}2}h_{\text{v}})$ where the $\rho_{\mathrm{l}1}$ and $\rho_{\mathrm{l}2}$ denote reference number densities per volume of the pure liquids 1 and 2, respectively.
@@
where $\mathbf{\underline{Q}}$ and $\mathbf{\underline{M}}$ are positive semi-definite and symmetric $6\times 6$ matrices representing the mobilities for the conserved and nonconserved parts of the dynamics, respectively. They are discussed in Appendix B. Note that the mobilities of Refs. [32, 33, 27, 6] are recovered as limiting cases. The functional $\mathcal{F}$ corresponds to the underlying free energy, i.e., the functional derivatives $\delta\mathcal{F}/\delta\psi_{i}$ represent chemical potentials. Their gradients within each phase and differences between phases drive the conserved (advection, diffusion) and nonconserved (evaporation/condensation, imbibition/desiccation) parts of the dynamics, respectively. We use
@@
where $\xi_{\mathrm{dg}}=\sqrt{1+|\nabla(h_{\text{d}}+h_{\text{b}})|^{2}}$ and $\xi_{\mathrm{bd}}=\sqrt{1+|\nabla h_{\text{b}}|^{2}}$ denote the metric factors for the two interfaces, and we use the $h$’s, $\zeta$’s and $\rho$’s as abbreviations for lengthy expressions only involving dependencies on the fields $\psi_{i}$. In total, there are bulk energies for the three phases, two interface energies and one wetting energy. The latter ultimately accounts for the brush-gas interface energy within the mesoscopic description [14, 6]. The specific forms of the individual contributions extend cases already considered the literature while recovering them as limiting cases. Examples include drops of simple nonvolatile and volatile liquids on polymer brushes [6], and drops of nonvolatile mixtures on rigid substrates [32]. Alternatively, instead of the thermodynamic formulation as gradient dynamics (1) one may present the same model in hydrodynamic form in terms of advective and diffusive fluxes and transfer rates between the phases. As the formulation is fully equivalent we abstain from showing it here, but see Ref. [6].
@@ 3 Results
Figure 2: Snapshots from a simulation of a spreading, evaporating and imbibing drop of a binary mixture of volatile partially wetting liquids on an initially dry polymer brush in 1d. On the left hand side an intermediate state is shown while the right hand side presents the final equilibrium state. Top, center and bottom panel give concentration/thickness profiles in gas, drop and brush, respectively. For initial state, parameters and a discussion of further details and main features see the main text and appendix C. Lateral boundary conditions are applied and only half the computational domain is shown.
@@ 4 Conclusion
In the present work, we have introduced a mesoscopic hydrodynamic model for drops/films of partially wetting volatile binary liquid mixtures on brush-covered substrates in a narrow-gap geometry. We have shown that the formulation as a gradient dynamics on an underlying free energy functional allows one to ’quasi-automatically’ account for all energetic cross-couplings between the individual fields and the resulting dynamic processes. In particular, the needed six scalar fields represent the two liquid components in the brush and the drop as well as the two corresponding vapors in the gas phase. All contributions to the energy have been detailed in Appendix A while the matrices of mobility functions for conserved and nonconserved parts of the dynamics are discussed in Appendix B. The energy accounts for concentration- and brush state-dependent capillarity and wettability, miscibility of the two components within drop and brush, as well as entropy in the gas phase. The developed model may account for the coupled spreading, absorption, imbibition, diffusion, evaporation/condensation, demixing/mixing and swelling dynamics that occurs when a drop of mixture is placed on a dry polymer brush. We have briefly illustrated this by employing an exemplary numerical time simulation of such a process. It has shown that the dynamic model seems to well describe the intricate coupled processes and may in the future serve in detailed studies of a variety of particular processes and phenomena related to the interaction of volatile mixtures and brushes. Examples include the experimental study of the influence of saturated vapors of different liquids on the wetting properties of brush-covered substrates that show atmosphere-dependent swelling [22]. Similarly, the presence of different organic vapors in the gas phase influences contact angle hysteresis for and resulting velocity of water drops on tilted brush-covered substrates as compared to the case without such vapors [40].
@@
However, the presented model only represents the first step in modeling the dynamics of droplets of volatile mixtures on adaptive substrates, a number of improvements are needed (see below). Here, we have shown that the gradient dynamics form of mesoscopic hydrodynamics is well suited to extend the reach of thin-film models towards adaptive substrates and liquid mixtures. Although, here we have only presented the full-curvature version of the model (indicated by the usage of the exact metric factors), a long-wave version is easily obtained by approximating the metric factors for small interface slopes, see, e.g., pertinent discussions in Refs. [12, 6]. We emphasize that the employed narrow-gap geometry allows one to study the influence of diffusion-limited evaporation in a thin-film setting. Here, we have extended the approach developed in Refs. [35, 27] toward a mixture of volatile liquids.
@@
However, there exist several options for future improvements of the presented basic model: (i) One may refine the brush energy accounting for more subtle interactions between brush, liquid mixture (and vapors). For instance, using the energy proposed in [25] should facilitate studies of the dynamics of various processes related to co-nonsolvency [41, 42, 43, 20, 22]. (ii) To alleviate the shortcoming of missing Marangoni-driven advective flows highly relevant for sessile drops of mixtures [44] or the merging of drops of different liquids [45] one may incorporate two additional fields describing the enrichment of one component at the brush-drop and the drop-gas interface, respectively. This is similar to the incorporation of a surfactant concentration field as done in the gradient dynamics formulations in Refs. [46, 34]. In principle, following the approach in section IV.B of Ref. [34] one may then apply an approximation to again eliminate the additional fields still keeping the Marangoni effect (at the price of slightly breaking the gradient structure). (iii) Beside the bulk energies one may also refine the concentration-dependencies of interface and wetting energies beyond the here-employed linear interpolations between known limiting cases (see, e.g., [47]). There, a discussion is needed that considers relations between the concentration dependencies of interface and wetting energies and the bulk mixture energies. (iv) Phase separation within the liquid mixture may result in micro-phase separation and very steep concentration gradients (as already visible in the contact line region in the studied example). To avoid such steep gradients one may incorporate an energetic penalty for concentration gradients, e.g., by using square-gradient terms in the energies for drop and brush phase similar to Cahn-Hilliard type energy densities [32]. (v) As in [6], we have assumed that transport of the two liquids within the brush is by diffusion only. This neglects any dynamic coupling between the lateral motion of the liquids within the brush and within the drop. Instead, one may model the motion within the brush similar to a (brush state-dependent) porous layer or, alternatively, introduce a brush state-dependent effective slip at the brush-drop interface (see, e.g., [48]). The former would add further nondiagonal terms to $\mathbf{\underline{Q}}$ while the latter would only amend existing nonzero components of $\mathbf{\underline{Q}}$.
@@
Note, finally, that the proposed model does not only allow to describe many dynamic processes for binary drops on adaptive substrates formed by polymer brushes. It may also serve as a blueprint for the development of similar models for the other types of adaptive substrates discussed in [3].
@@
none

### v2

@@ Abstract
We introduce a mesoscopic hydrodynamic model for drops of binary mixtures of volatile partially wetting liquids on brush-covered substrates, i.e., we model the coupled dynamics of spreading, evaporation, imbibition, diffusion and partial demixing of the two volatile components across the three phases – brush, drop and gas. The formulation of the model as gradient dynamics on an underlying free energy functional allows us to systematically account for cross-couplings between the six scalar fields needed to describe the dynamics of both components within each of the three phases. The energy accounts for concentration- and brush state-dependent capillarity and wettability, miscibility of the two components within drop and brush, and entropy in the gas. Finally, the usage of the model is illustrated by employing numerical time simulations to study the dynamics of a sessile drop.
@@ 1 Introduction
Overall, the statics and dynamics of the (de)wetting of simple nonvolatile liquids on rigid inert solid substrates are rather well understood [1]. However, at present, flexible and adaptive substrates, e.g., (hydro)gels and polymer brushes, are attracting much interest [2, 3, 4, 5, 6]. In general, films and drops of liquids on solid substrates are often modeled employing mesoscopic hydrodynamics in the form of thin-film (or lubrication) models that incorporate capillarity and wettability [7, 8, 9]. Particularly instructive and convenient is their formulation in the form of gradient dynamics on underlying energy functionals [10, 11, 12, 5, 6]. Then, consistency between macroscopic and mesoscopic modeling descriptions is ensured by relations between macroscopic interface energies and mesoscopic wetting energies [13, 14, 6].
@@
In the case of polymer brush-covered substrates [15, 16, 17, 18, 19, 20, 21, 22], thin-film models in gradient dynamics formulation [6] incorporate the free energy of the polymer brush, e.g., via an Alexander-de Gennes approach [23, 24]. In the simplest case, it accounts for mixing energy and entropy of the brush polymer and the imbibing liquid [25, 26]. As the polymers are grafted to the supporting solid, their translational entropy does not contribute and is replaced by entropic stretching. The resulting thin-film models may be employed to investigate the spreading of drops of nonvolatile [26, 6] and volatile [27, 6] liquids on polymer brushes, the stick-slip motion of three-phase contact lines advancing over an initially dry brush [28, 6], as well as sliding drops on such substrates [6]. Thereby, evaporation and vapor dynamics are incorporated into the gradient dynamics model following an approach recently proposed for drops of volatile simple one-component liquids in a gap [29].
Here, we provide a major extension of such models by incorporating the description of liquid mixtures, i.e., for a binary mixture both components are present in brush, drop and vapor. Although, this is of large present interest [25, 30, 19, 22], to our knowledge, the coupled dynamics of spreading, evaporation, imbibition, diffusion and (partial) demixing of the two volatile components across the three phases – brush, drop and gas – has not yet been modeled. We emphasize that the gradient dynamics-based approach “semi-automatically” guarantees thermodynamic consistency and nevertheless incorporates the relevant dissipation channels. Otherwise, this consistency might easily get lost when modeling such complex interface-dominated physico-chemical systems. Note, however, that we restrict our attention to isothermal settings as gradient dynamics-based thin-film models that incorporate heat transport are, to our knowledge, not yet available.
Note, however, binary mixtures of nonvolatile liquids on rigid inert solids are already described by mesoscopic hydrodynamics models [31, 32, 33, 34]. This allows, e.g., to investigate dewetting processes that are triggered by the coupling of fluctuations in film thickness and concentration [33]. Surface activity due to insoluble or soluble surfactants may also be incorporated [35, 14].
@@
As mentioned above, evaporation and vapor diffusion for volatile liquids are accounted for in thin-film modeling by confining the system into a small gap between two parallel plates as then vapor diffusion is predominantly lateral while concentrations equilibrate fast across the gap [29]. In contrast to earlier thin-film models including evaporation, see Refs. [7, 8] and the discussion and references in the introduction of Ref. [29], this allows one to describe the full parameter range from phase transition-limited to diffusion-limited evaporation/condensation. An application of the evaporation model of Ref. [29] to droplets of a simple weakly volatile liquid spreading on a brush [27] shows that even a small lateral vapor leakage from a closed system may result in the formation of a long-lived stationary halo of macroscopic extension in the brush profile.
@@
The model is introduced in section 2 by detailing the six scalar fields representing the amount of the two substances in the three phases brush, drop and gas, and the corresponding dynamic equations. Details regarding the underlying energy functional and mobility functions are provided in Appendices A and B, respectively. Using the established model, section 3 presents exemplary results obtained by numerical time simulations. Finally, section 4 summarizes our results, points out possible applications of the developed model, and proposes future extensions that may alleviate present limitations.
@@ 2 Mesoscopic hydrodynamic model in gradient dynamics form
We consider an isothermal system as sketched in Fig. 1: A drop of a binary mixture of volatile liquids is located on a swelling polymer brush. Described are the amounts of the two liquid components in the drop and brush by effective heights $\zeta_{\mathrm{d}i}$ and $\zeta_{\mathrm{b}i}$ (with $i=1,2$), respectively, the thickness of the drop is $h_{\text{d}}=\zeta_{\text{d1}}+\zeta_{\text{d2}}$ while the brush thickness is $h_{\text{b}}=\zeta_{\text{b1}}+\zeta_{\text{b2}}+\zeta_{\text{p}}$. Here, $\zeta_{\text{p}}$ is the uniform and constant effective height of polymers, i.e., the dry brush height $\zeta_{\text{p}}=\sigma Nl_{\text{K}}$, with $\sigma$ the relative grafting density, $N$ the number of monomers per polymer chain and $l_{\text{K}}$ the Kuhn length. We consider the system to be positioned inside a narrow gap between two horizontal plates of distance $d$, therefore, the height of the gas phase is $h_{\text{v}}=d-h_{\text{d}}-h_{\text{b}}$.
Thin-film models for liquid films and drops involving mixtures are conveniently written in gradient dynamics form employing (effective) heights [33, 34, 35, 36]. However, when vapor is incorporated, it is preferable to express the amounts of each component in each phase - brush, drop and gas - as particle number densities per substrate area $\psi_{i}(\mathbf{r},t)$, where $\mathbf{r}=(x,y)^{T}$ denotes the substrate coordinates [27, 6]. Here, we use the six fields $(\psi_{1},\psi_{2},\psi_{3},\psi_{4},\psi_{5},\psi_{6})=(\rho_{\mathrm{l}1}\zeta_{\text{d1}},\rho_{\mathrm{l}1}\zeta_{\text{b1}},\rho_{\mathrm{v}1}h_{\text{v}},\rho_{\mathrm{l}2}\zeta_{\text{d2}},\rho_{\mathrm{l}2}\zeta_{\text{b2}},\rho_{\mathrm{v}2}h_{\text{v}})$ where the $\rho_{\mathrm{l}1}$ and $\rho_{\mathrm{l}2}$ denote reference number densities per volume of the pure liquids 1 and 2, respectively. We emphasize that the system geometry is identical to the one presented in [6, 27]. However, here we provide a major extension of earlier work by incorporating the description of a binary mixture, i.e., two substances (instead of only one) contribute to the dynamics in each bulk phase.
@@
where $\mathbf{\underline{Q}}$ and $\mathbf{\underline{M}}$ are positive semi-definite and symmetric $6\times 6$ matrices representing the mobilities for the conserved and nonconserved parts of the dynamics, respectively. The individual components of both mobility matrices are discussed in Appendix B. Note that the mobilities of Refs. [33, 34, 27, 6] are recovered as limiting cases. The functional $\mathcal{F}$ corresponds to the underlying free energy, i.e., the functional derivatives $\delta\mathcal{F}/\delta\psi_{i}$ represent chemical potentials. Their gradients within each phase and differences between phases drive the conserved (advection, diffusion) and nonconserved (evaporation/condensation, imbibition/desiccation) parts of the dynamics, respectively. We use
@@
where $\xi_{\mathrm{dg}}=\sqrt{1+|\nabla(h_{\text{d}}+h_{\text{b}})|^{2}}$ and $\xi_{\mathrm{bd}}=\sqrt{1+|\nabla h_{\text{b}}|^{2}}$ denote the metric factors for the two interfaces, and we use the $h$’s, $\zeta$’s and $\rho$’s as abbreviations for lengthy expressions only involving dependencies on the fields $\psi_{i}$. In total, there are bulk energies for the three phases, two interface energies and one wetting energy. The latter ultimately accounts for the brush-gas interface energy within the mesoscopic description [14, 6]. The specific forms of the individual contributions extend cases already considered the literature while recovering them as limiting cases. Examples include drops of simple nonvolatile and volatile liquids on polymer brushes [6], and drops of nonvolatile mixtures on rigid substrates [33]. The individual expressions of all contributions to $\mathcal{F}$ are discussed in Appendix A.
Alternatively, instead of the thermodynamic formulation as gradient dynamics (1), one may present the same model in hydrodynamic form in terms of advective and diffusive fluxes and transfer rates between the phases. As the formulation is fully equivalent we abstain from showing it here, but see Ref. [6] for details regarding the “translation” between the two formulations (at the example of a simple liquid).
@@ 3 Results
Figure 2: Snapshots from a simulation of a spreading, evaporating and imbibing drop of a binary mixture of volatile partially wetting liquids on an initially dry polymer brush in 1d. On the left hand side an intermediate state is shown while the right hand side presents the final equilibrium state. Top, center and bottom panels give concentration/thickness profiles in gas, drop and brush, respectively. For initial state, parameters, and a discussion of further details and main features see the main text and appendix C. Lateral boundary conditions are applied and only half the computational domain is shown.
@@ 4 Conclusion
We have introduced a mesoscopic hydrodynamic model for drops of a binary mixture of volatile liquids on brush-covered substrates in a narrow-gap geometry. Our work corresponds to a major extension of earlier work on gradient dynamics-based thin-film descriptions of drops of volatile simple liquids in a gap [29] as well as of nonvolatile and volatile simple liquids on brush-covered substrates [6]. Although, here we have only presented the full-curvature version of the model (indicated by the usage of the exact metric factors), a long-wave version is easily obtained by approximating the metric factors for small interface slopes, see, e.g., pertinent discussions in Refs. [12, 6].
We have shown that the formulation as a gradient dynamics on an underlying free energy functional allows one to ’quasi-automatically’ account for all energetic cross-couplings between the individual fields and the resulting dynamic processes. In particular, the needed six scalar fields represent the two liquid components of the mixture within the brush and the drop as well as the two components of the vapor in the gas phase. All contributions to the energy have been detailed in Appendix A while the matrices of mobility functions for the conserved and nonconserved parts of the dynamics are discussed in Appendix B. The energy accounts for concentration- and brush state-dependent capillarity and wettability, miscibility of the two components within drop and brush, as well as entropy in the gas phase. The developed model may account for the coupled spreading, absorption, imbibition, diffusion, evaporation/condensation, demixing/mixing and swelling dynamics that occurs when a drop of binary mixture is placed on a dry polymer brush. We have briefly illustrated this by employing an exemplary numerical time simulation of such a process. It has shown that the dynamic model seems to well describe the intricate coupled processes and may in the future serve in detailed studies of a variety of particular processes and phenomena related to the interaction of volatile mixtures and brushes. Examples include the experimental study of the influence of saturated vapors of different liquids on the wetting properties of brush-covered substrates that show atmosphere-dependent swelling [22]. Similarly, the presence of different organic vapors in the gas phase influences contact angle hysteresis for and resulting velocity of water drops on tilted brush-covered substrates as compared to the case without such vapors [40].
@@
Note, however, that the developed model only represents a first step in modeling the dynamics of droplets of mixtures of volatile liquids on adaptive substrates. Although, here we have shown that the gradient dynamics form of mesoscopic hydrodynamics is well suited to extend the reach of thin-film models toward liquid mixtures, a number of improvements of the presented basic model should be pursued in the future: (i) One may refine the brush energy accounting for more subtle interactions between brush, liquid mixture (and vapors). For instance, using the energy proposed in [25] should facilitate studies of the dynamics of various processes related to co-nonsolvency [41, 42, 43, 20, 22]. (ii) The present model does not account for advective flows driven by solutal Marangoni forces resulting from concentration dependencies of interface energies. However, they are highly relevant for sessile drops of mixtures [44] or the merging of drops of different liquids [45]. They may be incorporated into the presented model by employing two additional fields that describe the enrichment of components at the brush-drop and the drop-gas interface, respectively. This is similar to the incorporation of a surfactant concentration field as done in the gradient dynamics formulations in Refs. [46, 35]. In principle, following the approach in section IV.B of Ref. [35] one may then apply an approximation to again eliminate the additional fields still keeping the Marangoni effect (at the price of slightly breaking the gradient structure). (iii) Beside the bulk energies one may also refine the concentration-dependencies of interface and wetting energies beyond the here-employed linear interpolations between known limiting cases (see, e.g., [47]). There, a discussion is needed that considers relations between the concentration dependencies of interface and wetting energies and the bulk mixture energies. (iv) Phase separation within the liquid mixture may result in micro-phase separation and very steep concentration gradients (as already visible in the contact line region in the studied example). To avoid such steep gradients one may incorporate an energetic penalty for concentration gradients, e.g., by using square-gradient terms in the energies for drop and brush phase similar to Cahn-Hilliard type energy densities [33]. (v) As in [6], we have assumed that transport of the two liquids within the brush is by diffusion only. This neglects any dynamic coupling between the lateral motion of the liquids within the brush and within the drop. Instead, one may model the motion within the brush similar to a (brush state-dependent) porous layer or, alternatively, introduce a brush state-dependent effective slip at the brush-drop interface (see, e.g., [48]). The former would add further nondiagonal terms to $\mathbf{\underline{Q}}$ while the latter would only amend existing nonzero components of $\mathbf{\underline{Q}}$. (vi) Our work has assumed an isothermal setting, although effects driven by latent heat and subsequent diffusion of heat are often relevant when considering evaporation [49]. To alleviate this limitation one would need to consistently incorporate the transport of heat into the gradient dynamics formalism. To our knowledge, this has not yet been done.
@@
Note, finally, that the proposed model does not only allow one to describe dynamic processes for binary drops on adaptive substrates formed by polymer brushes. It may also serve as a blueprint for the development of similar models for most other types of adaptive substrates discussed in [3].
@@
This version of the article has been accepted for publication, after peer review but is not the Version of Record and does not reflect post-acceptance improvements, or any corrections. The Version of Record is available online at: http://dx.doi.org/10.1140/epjs/s11734-024-01169-4.
@@
Data Availability Statement
The underlying data needed to reproduce shown results are publicly available at the data repository zenodo (10.5281/zenodo.10696697).